OpenStudy (anonymous):

Given two points on the line, input the equation of the line in standard form Ax + By = C. Reduce all fractional answers to lowest terms. (7, -3), (4, -8)

OpenStudy (anonymous):

start by finding slope m=(y2-y1)/(x2-x1)

OpenStudy (anonymous):

y2=-8 y1=-3 x2=4 x1=7

OpenStudy (anonymous):

m=(-8--3)/(4-7)=(-5/-3)=(5/3)

OpenStudy (anonymous):

now that you have m plug one of the points and m into the point slope equation y-y1=m(x-x1)

OpenStudy (anonymous):

you gotta take the last step here. can't answer them all for ya